OpenStudy (anonymous):

find the general term an for the geometric sequence. a1 = -5 and a2 = 20

OpenStudy (anonymous):

Note that, for a geometric series: \[ \frac{a_{n+1}}{a_{n}}=k \]For some \(k\). Hence: \[ \frac{20}{-5}=-4=k \]And, our first term \(a_1=-5\), so we have: \[ a_n=a_1(k)^{n-1}=(-5)(-4)^{n-1} \]

OpenStudy (anonymous):

Ok....so is the answer 20^n-1

OpenStudy (anonymous):

Careful, mate, since you can't multiply exponent bases.

OpenStudy (anonymous):

ok...so an = (-5)(-4)^n-1